# GIGA: Generalizable Sparse Image-driven Gaussian Avatars

### v1

@@ Abstract
Driving a high-quality and photorealistic full-body human avatar, from only a few RGB cameras, is a challenging problem that has become increasingly relevant with emerging virtual reality technologies. To democratize such technology, a promising solution may be a generalizable method that takes sparse multi-view images of an unseen person and then generates photoreal free-view renderings of such identity. However, the current state of the art is not scalable to very large datasets and, thus, lacks in diversity and photorealism. To address this problem, we propose a novel, generalizable full-body model for rendering photoreal humans in free viewpoint, as driven by sparse multi-view video. For the first time in literature, our model can scale up training to thousands of subjects while maintaining high photorealism. At the core, we introduce a MultiHeadUNet architecture, which takes sparse multi-view images in texture space as input and predicts Gaussian primitives represented as 2D texels on top of a human body mesh. Importantly, we represent sparse-view image information, body shape, and the Gaussian parameters in 2D so that we can design a deep and scalable architecture entirely based on 2D convolutions and attention mechanisms. At test time, our method synthesizes an articulated 3D Gaussian-based avatar from as few as four input views and a tracked body template for unseen identities. Our method excels over prior works by a significant margin in terms of cross-subject generalization capability as well as photorealism. The project page is available at vcai.mpi-inf.mpg.de/projects/GIGA.
@@ 1 Introduction
Driving your own full-body and photoreal virtual avatar from a scarce and affordable multi-view (i.e. four), camera setups has the potential to revolutionize communication, gaming, and remote collaboration. However, there are to date modeling challenges that remain unsolved: 1) Achieving photorealism and fidelity despite sensor scarcity and limited input data. 2) True generalization to unseen, novel identities. In this work, we attempt to jointly solve them by leveraging recent large-scale data capture efforts. Notably, this requires a generalizable method to synthesize digital humans at test time in a simple feed-forward manner, which is the subject of this work.
@@
Recent related works have focused on person-specific avatars, i.e., a learned representation that is trained per subject on dense dome-like camera setups. These representations may involve meshes [5, 1, 54], neural radiance fields [33, 11, 48], points [45], or volumetric primitives [36] such as 3D Gaussian splats [62, 31, 28]. Democratizing such technology is difficult since, for high-quality results, a dense camera dome is required before the character can be driven at inference time. Some methods aim to utilize simpler capture setups, e.g., monocular images [56, 57] or videos [49, 8]. However, due to the much scarcer input, their visual quality often falls short compared to multi-view methods. Single-image to 3D reconstruction methods [40, 41, 56, 57, 10] also primarily focus on adequate geometry reconstruction and produce a 3D asset, drivable with skeletal motion only, which typically lacks the skeletal motion-dependent geometry and appearance changes.
@@
Designing a high-quality approach for lightweight human reconstruction and rendering, thus, remains an open problem: the model must accurately represent diverse human appearances, body types, and clothing configurations and it must correctly derive pose-dependent appearance changes from the scarce input signal. While previous works [15, 30, 51, 16] have shown promising first steps towards identity generalization, those methods were trained and evaluated on small-scale datasets [34] with few training subjects. Moreover, the quality and efficiency of these works are constrained by the limitations of implicit neural representations and non-scalable network architectures. More recently, large-scale datasets [3, 55] helped to pave the way for truly generalizable human avatar methods. However, training and evaluating such models at scale demands particular consideration. First, a generalizable model has to learn meaningful feature representations to scale, both, to the training set and outside of it. Second, the model architecture must be computationally and memory-efficient to capture fine detail and enable large-scale training. Third, at the core of such a method, the 3D representation has to yield high-quality reconstruction with minimal rendering time to maintain fast backpropagation during training and to ensure fast performance at test time.
@@
To address these challenges, we propose GIGA, a feed-forward method to synthesize personalized virtual human avatars from sparse input views and a tracked body template at inference. Notably, for unseen subjects, GIGA requires no personalized training on dense dome data. At the core, we project sparse-view image information into the UV space of the SMPL-X model [32] while the digital human predicted by GIGA is represented as a texel-aligned Gaussian avatar [16, 12] greatly simplifying the task to a 2D-to-2D image translation. We propose a MultiHeadUNet, a UNet with multiple encoding and decoding heads, which takes the texture containing the projected image information as well as shape and motion codes and regresses per-texel Gaussian appearance and geometry parameters. In detail, we employ cross attention to inject motion information in the model, and skip-connections to propagate learning signal at different spatial scales inside of the network. Our architecture design choices ensure reliable learning of intrinsic feature statistics from the training data while maintaining person-specific information contained in the model inputs. The predicted Gaussian parameters can be posed into 3D space using the respective SMPL-X body pose, which allows rendering 2D images that can be compared to the ground truth during training. In summary, our contributions are:
@@
  • We introduce GIGA, a novel generalizable human avatar reconstruction and rendering approach operating in a 4-view setup in a feed-forward regime at test time.
@@
  • We design GIGA as a MultiHeadUNet with motion conditioning relying on attention mechanism to synthesize high-quality 3D Gaussian-based avatars with dynamic appearance changes.
  • GIGA for the first time in literature can be effectively trained on thousands of human identities, and shows compelling results in large-scale generalization.
@@
We demonstrate the generalization capabilities of GIGA through a comprehensive evaluation on large-scale datasets MVHumanNet [55], DNA-Rendering [3], and THuman2.0 [59]. Experimental results show that GIGA significantly outperforms prior works in visual quality as well as identity and pose generalization.
@@ 2 Related Work
Person-specific Human Capture and Rendering. Neural implicit approaches for novel view synthesis [27, 47, 29] were primarily designed for per-scene optimization. This paradigm naturally extended to human modeling through various implicit representations: radiance fields [34, 44, 11], signed distance functions [48, 52, 63], and occupancy fields [40, 41, 26]. These methods typically rely on parametric human body models [24, 32] as geometric proxies to initialize neural implicit representations. The inherent limitations of generic parametric templates in capturing person-specific geometry led to methods utilizing personalized mesh templates [20, 5, 36], which demonstrate superior reconstruction quality, mimicking pose-dependent appearance changes. Drivable Volumetric Avatars (DVA) [36] target telepresence applications by representing human avatars as mixtures of volumetric primitives [23], regressed from texel-aligned image and pose features extracted from 3 input views. Holoported Characters [42] combines a personalized mesh template with dynamic feature textures predicted from partial texture and normal maps, operating with 4 input views. This method achieves 4K resolution rendering through a superresolution network applied to rendered feature images. 3D Gaussian Splatting [13] marked a significant advancement in, both, rendering quality and computational efficiency. 3D Gaussians have been successfully adapted for free-viewpoint human avatar rendering in multi-view [31, 28, 12, 65] and monocular [7, 8, 35] setups. These approaches remain constrained to person-specific optimization and cannot be easily extended to cross-identity training.
@@
Figure 2: Method Overview. GIGA generates dynamic textures of 3D Gaussians for photoreal avatars from 4 input views $\mathcal{I}_{k}$ and a tracked body template $(\boldsymbol{\theta},\boldsymbol{\beta})$. An initial RGB texture $\mathbf{T}_{\mathrm{uv}}$ is gathered from the input images and passed to the appearance encoder $\mathcal{E}_{\mathrm{a}}$ to extract appearance features $\mathbf{F}^{\mathrm{a}}_{\mathrm{uv}}$. A canonical position map $\mathbf{T}_{\mathbf{x}_{0}}$ is processed by the geometry encoder $\mathcal{E}_{\mathrm{g}}$ into geometry features $\mathbf{F}^{\mathrm{g}}_{\mathrm{uv}}$. Both intermediate features are motion-dependent via cross-attention conditioning on the observed pose $\boldsymbol{\theta}$. Gaussian texel maps are regressed by separate decoders, each for an individual group of parameters. Skip-connections (dashed lines) propagate intermediate features from encoders to decoders. Output of GIGA is articulated with linear blend skinning.
@@
Generalizable Human Capture and Rendering. Recent research has addressed the challenge of synthesizing free-viewpoint videos of human performances from sparse multi-view captures with generalization across subjects [6, 64, 9]. Neural Human Performer (NHP) [15] integrates pixel-aligned visual features with skeletal pose information extracted from multi-view video sequences through cross-attention. However, NHP’s performance degrades in regions occluded from input viewpoints due to its direct operation in the observed pose space. TransHuman [30] addresses this limitation by learning to synthesize a NeRF-based representation in the canonical pose space, employing a transformer architecture that processes individual body parts as tokens. While NHP utilizes cross-attention to fuse temporal features first and frame-specific multi-view features next, TransHuman constructs an implicit representation through N-nearest neighbors’ token aggregation. Neural Novel Actor (NNA) [51] disentangles appearance and geometry by processing spatial point features and SMPL surface features through a graph CNN, enabling separate prediction of person-specific appearance and pose-dependent deformation, as proposed by Liu et al. [21].
The identity generalization capabilities of these methods remain limited due to training on datasets with at most 10 subjects. Additionally, their reliance on implicit representations and network architecture design imposes significant computational costs during rendering, further increasing training time on larger datasets. In contrast, Generalizable Human Gaussians (GHG) [16] introduces an alternative approach for reconstructing static human models from sparse input views. By leveraging a large-scale dataset of textured human 3D scans [59], GHG demonstrates generalization to unseen subjects while achieving near real-time performance by representing avatars as multiple scaffolds of 3D Gaussians in the UV space. However, since GHG is supervised with ground-truth 3D data, its training cannot be easily extended to multi-view datasets.
@@ 3 Method
GIGA aims at mapping sparse image observations (4 views) of an unseen human to a 3D Gaussian-based avatar that can be rendered from a free viewpoint. This is achieved while relying only on an estimated parametric body model that captures the rough shape and pose of the person. Importantly, our method requires no personalized training.
This section first formalizes the problem setting and relevant background knowledge (Sec. 3.1). It then introduces our generalizable human representation (Sec. 3.2), our training strategy (Sec. 3.3), and implementation details (Sec. 3.4). Fig. 2 provides an overview.
@@ 3.1 Problem Setting and Background
Problem Setting. GIGA assumes a collection of multi-view videos of several subjects with per-frame subject segmentation masks for training. Each subject is captured by $\hat{K}=16$ calibrated cameras, with $\pi_{\hat{k}}$ denoting the projection matrix for camera $\hat{k}$. Each video frame is annotated with SMPL-X [32] parameters released as additional labels in most multi-view human performance capture datasets  [55, 3, 59]. At test time, GIGA takes as input sparse-view videos as driving signals and regresses 3D Gaussian splats that faithfully preserve the appearance and clothing dynamics from these driving videos, even for unseen identities. For training, we select $K=4$ sparse camera views and use native SMPL-X parameters to construct the input for the encoder network. The images from the remaining $K^{\prime}=12$ views serve as ground truth to supervise the outputs decoded by the network.
@@ 3.2 Generalizable Human Representation
Figure 4: Qualitative Comparison against Static Method. While sharing similar concepts with GHG [16], GIGA (Ours) outperforms it on THuman2.0 [59] dataset thanks to our design choices. GHG also cannot be straightforwardly applied to dynamic sequences, which is not the case for GIGA.
The original 3D Gaussian representation is a point cloud where each primitive is essentially independent from the others. A naive approach would optimize separate sets of 3D Gaussians for each training identity. Within such a set, each Gaussian is optimized almost independently from every other, leading to an absence of shared statistical information between neighboring Gaussians. Moreover, it would be hard to establish semantic correspondence between different sets of Gaussians.
Inspired by previous personalized, animatable Gaussian avatar methods [31, 12, 50], we represent virtual humans as texel-aligned 3D Gaussian maps within the texture space $\mathcal{M}_{\mathrm{uv}}$ of the SMPL-X template mesh. Anchoring 3D Gaussians to the texture space $\mathcal{M}_{\mathrm{uv}}$ solves the aforementioned problems: texel positions are fixed on the template surface, making 3D Gaussians assigned to same texels semantically similar across various characters; 2D textures can be efficiently processed by convolutional neural networks; and Gaussians from the same region of the texture space will be entangled thanks to the locality bias of convolutions. As such, GIGA can learn cross-subject information and predict the final texel-aligned Gaussian avatar with a single pass of a 2D convolutional neural network.
Input Appearance Encoding. We use information from the input views to capture identity-specific appearance with pose-dependent variations. To transform this information from image to texture space, we first compute partial textures $\mathbf{T}_{\mathrm{uv},k}$ for each view $\pi_{k}$ by projecting each pixel onto the 3D surface of SMPL-X. This 3D point can be mapped into texture space using the UV mapping of SMPL-X such that the final image pixel color is projected onto the respective 2D location in the texture map. We then fuse those textures according to visibility resulting in the final texture $\mathbf{T}_{\mathrm{uv}}\in\mathbb{R}^{T\times T\times 3},T=512$ in the UV space $\mathcal{M}_{\mathrm{uv}}$.
@@
which extracts appearance features $\mathbf{F}_{\mathrm{uv}}^{a}\in\mathbb{R}^{T_{f}\times T_{f}\times d}$, that encode character-specific appearance and identity information. The encoder $\mathcal{E}_{\mathrm{a}}$ consist of 2D convolutional downsampling residual blocks. Feature maps $\mathbf{H}^{\mathrm{a}}$ from also taken from the downsampling levels for later usage in decoding.
$\mathcal{E}_{\mathrm{a}}$ is conditioned on the motion embedding $\mathbf{y}_{\mathrm{m}}$. Following Rombach et al. [37], this conditioning is implemented using cross-attention layers as final layers of the encoder.
Motion Embeddings. The input texture $\mathbf{T}_{\mathrm{uv}}$ contains only a localized pose-dependent learning signal from the template surface. We additionally use an MLP-based motion encoder $\mathcal{E}_{\mathrm{m}}$ to construct motion embeddings for SMPL-X poses $\boldsymbol{\theta}$, which adds global body motion awareness at the encoding stage:
@@
Gaussian Primitives Regression. For the decoding stage, we design three separate decoders (Fig. 2): $\mathcal{D}_{\mathrm{a}}$ for the appearance, $\mathcal{D}_{\mathrm{p}}$ for Gaussian scales, quaternions and opacities, and $\mathcal{D}_{\mathrm{g}}$ for Gaussian offsets. All three share the same convolutional architecture and receive appearance and geometry features $\mathbf{F}_{\mathrm{uv}}^{\mathrm{a}},\mathbf{F}_{\mathrm{uv}}^{\mathrm{g}}$ as inputs:
@@
The output map $\mathcal{G}^{\prime}_{\mathrm{uv}}$ contains RGB color channels $\mathbf{c}$, quaternions $\mathbf{q}_{0}$, normalized scales $\mathbf{s^{\prime}}$, opacities $\alpha$, and offsets $\delta\mathbf{x}_{0}$, which are defined in the canonical T-pose. To utilize the representational power of the shared texel space and to propagate semantic information from encoders to decoders at different spatial scales, we use UNet-like [39] skip-connections, building a feature propagation bridge from each encoder to each decoder. We stack intermediate feature maps $\mathbf{H}^{\mathrm{a}}$ and $\mathbf{H}^{\mathrm{g}}$ along the feature dimension and feed them to every decoder at each corresponding upsampling layer. To convert a texel-aligned Gaussian map to the observed pose space $\boldsymbol{\theta}$, we use linear blend skinning applied to both offsets $\delta\mathbf{x}_{0}$ and quaternions $\mathbf{q}_{0}$: $\delta\mathbf{x}=\operatorname{lbs}\left(\delta\mathbf{x}_{0},\boldsymbol{\theta}\right);\mathbf{q}=\operatorname{lbs}\left(\mathbf{q}_{0},\boldsymbol{\theta}\right)$. We also multiply normalized Gaussian scales $\mathbf{s}^{\prime}$ by a hyperparameter $\rho$: $\mathbf{s}=\rho\mathbf{s}^{\prime}$, value $\rho=5e^{-3}$ was empirically found adequate to provide sufficient coverage of the template with Gaussians.

### v2

# GIGA: Generalizable Sparse Image-driven Gaussian Humans

@@ Abstract
Driving a high-quality and photorealistic full-body virtual human from a few RGB cameras is a challenging problem that has become increasingly relevant with emerging virtual reality technologies. A promising solution to democratize such technology would be a generalizable method that takes sparse multi-view images of any person and then generates photoreal free-view renderings of them. However, the state-of-the-art approaches are not scalable to very large datasets and, thus, lack diversity and photorealism. To address this problem, we propose GIGA, a novel, generalizable full-body model for rendering photoreal humans in free viewpoint, driven by a single-view or sparse multi-view video. Notably, GIGA can scale training to a few thousand subjects while maintaining high photorealism and synthesizing dynamic appearance. At the core, we introduce a MultiHeadUNet architecture, which takes an approximate RGB texture accumulated from a single or multiple sparse views and predicts 3D Gaussian primitives represented as 2D texels on top of a human body mesh. At test time, our method performs novel view synthesis of a virtual 3D Gaussian-based human from 1 to 4 input views and a tracked body template for unseen identities. Our method excels over prior works by a significant margin in terms of identity generalization capability and photorealism.
@@ 1 Introduction
Driving your own full-body and photoreal virtual double from a scarce and affordable single or sparse-view (up to 4) camera setup has the potential to revolutionize communication, gaming, and remote collaboration. However, there are to date modeling challenges that remain unsolved: 1) Achieving photorealism and fidelity despite sensor scarcity and limited input data. 2) Generalization ability to novel identities unseen during training. In this work, we attempt to jointly solve them by leveraging recent large-scale data capture efforts. Notably, this requires a generalizable method to synthesize digital humans at test time in a simple feed-forward manner, which is the subject of this work.
@@
Recent related works have most focused on person-specific avatars, i.e., a learned representation that is trained per subject on dense dome-like camera setups, which can be later articulated with arbitrary skeletal motion. These representations may involve meshes [6, 1, 60], neural radiance fields [38, 12, 54, 70], points [51], or volumetric primitives [41] such as 3D Gaussians [23, 35, 32]. Democratizing such technology is difficult since, for high-quality results, a dense camera dome is required before the character can be driven at inference time. Some methods aim to utilize simpler capture setups, e.g., monocular images [62, 63] or videos [55, 9]. Nevertheless, due to the much scarcer input, their visual quality often falls short compared to multi-view methods. Single-image to 3D reconstruction methods [45, 46, 62, 63, 11, 72, 74] also primarily focus on adequate geometry reconstruction and produce a 3D asset, drivable with skeletal motion only, which typically lacks the skeletal motion-dependent geometry and appearance changes.
@@
Designing a high-quality approach for lightweight human reconstruction and rendering, thus, remains an open problem: the model must accurately represent diverse human appearances, body types, and clothing configurations, and it must correctly derive pose-dependent appearance changes from the scarce input signal. While previous works [18, 34, 57, 19] have shown promising first steps towards identity generalization, those methods were trained and evaluated on small-scale datasets [39] with few training subjects. Moreover, the quality and efficiency of these works are constrained by the limitations of implicit neural representations and non-scalable network architectures. More recently, large-scale datasets [3, 61] helped to pave the way for truly generalizable human modeling methods. As was shown by Guo et al. [5], learning from a large data collection enables construction of a rich human avatar prior, albeit requiring preparation of personalized body templates beforehand. Nonetheless, training and evaluating such models at scale demands particular consideration: First, a generalizable model has to learn meaningful feature representations to scale, both to the training set and outside of it. Second, the model architecture must be computationally and memory-efficient to capture fine detail and enable large-scale training. Third, the underlying 3D representation has to yield high-quality, efficient rendering.
@@
To address these challenges, we propose GIGA, a feed-forward method to synthesize virtual 3D-Gaussian-based humans from single or sparse input views and a tracked body template at inference for free-view rendering. Notably, for unseen subjects, GIGA requires no personalized training on dense dome data and no fitting of personalized mesh templates. At the core, we project sparse-view image features into the UV space of the SMPL-X model [36] while the virtual human predicted by GIGA is represented as a set of texel-aligned 3D Gaussians [19, 14, 72, 64]. We propose a MultiHeadUNet, a UNet with multiple encoding and decoding heads, which takes the texture containing the projected image information as well as shape and motion codes and regresses per-texel Gaussian appearance and geometry parameters. In detail, we employ cross attention to inject motion information into the model, and skip-connections to propagate the learning signal at different spatial scales within the network. Our architecture design choices ensure reliable learning of intrinsic feature statistics from the training data while maintaining person-specific information contained in the model inputs. The predicted 3D Gaussians can be posed into 3D space using the respective SMPL-X body pose and further rendered onto novel views. In summary, our contributions are:
@@
  • We introduce GIGA, a novel generalizable method for dynamic novel view synthesis of virtual humans operating in a single or sparse view setup in a feed-forward regime at test time.
@@
  • We design a MultiHeadUNet with motion conditioning to synthesize high-quality texel-aligned 3D Gaussians with dynamic appearance changes.
  • We train GIGA on a large-scale collection of multi-view videos, after which it generalizes to unseen identities within and across datasets, producing dynamic novel view renders even with as few as a single input video.
@@
We demonstrate the generalization capabilities of GIGA through a comprehensive evaluation on large-scale multi-view [61, 3, 66] and monocular videos [13] datasets. Experimental results show that GIGA significantly outperforms prior works in visual quality as well as identity and pose generalization.
@@ 2 Related Work
Person-specific Human Capture and Rendering. Neural implicit approaches for novel view synthesis [31, 53, 33] were primarily designed for per-scene optimization. This paradigm naturally extended to human modeling through various implicit representations: radiance fields [39, 50, 12], signed distance functions [54, 58, 70], and occupancy fields [45, 46, 30]. These methods typically rely on parametric human body models [28, 36] as geometric proxies to initialize neural implicit representations. The inherent limitations of generic parametric templates in capturing person-specific geometry led to methods utilizing personalized mesh templates [24, 6, 41], which demonstrate superior reconstruction quality, mimicking pose-dependent appearance changes. Drivable Volumetric Avatars (DVA) [41] target telepresence applications by representing human avatars as mixtures of volumetric primitives [27], regressed from texel-aligned image and pose features extracted from 3 input views. Holoported Characters [48] combines a personalized mesh template with dynamic feature textures predicted from partial texture and normal maps, operating with 4 input views, which yields 4K rendering resolution. 3D Gaussian Splatting [15] marked a significant advancement in both rendering quality and computational efficiency. 3D Gaussians have been successfully adapted for free-viewpoint human avatar rendering in multi-view [35, 32, 14, 73] and monocular [8, 9, 40] setups, mostly constrained to person-specific optimization without cross-identity training.
@@
Figure 2: Method Overview. GIGA generates dynamic textures of 3D Gaussians for photoreal avatars from sparse input views $\mathcal{I}_{k}$ and a tracked body template $(\boldsymbol{\theta},\boldsymbol{\beta})$. An initial RGB texture $\mathbf{T}_{\mathrm{uv}}$ is gathered from the input images and passed to the appearance encoder $\mathcal{E}_{\mathrm{a}}$ to extract appearance features $\mathbf{F}^{\mathrm{a}}_{\mathrm{uv}}$. A canonical position map $\mathbf{T}_{\mathbf{x}_{0}}$ is processed by the geometry encoder $\mathcal{E}_{\mathrm{g}}$ into geometry features $\mathbf{F}^{\mathrm{g}}_{\mathrm{uv}}$. Both intermediate features are motion-dependent via cross-attention conditioning on the observed pose $\boldsymbol{\theta}$. Gaussian texel maps are regressed by separate decoders, each for an individual group of parameters. Skip-connections (dashed lines) propagate intermediate features from encoders to decoders. Output of GIGA is articulated with linear blend skinning.
@@
Generalizable Human Capture and Rendering. Recent research has addressed the challenge of synthesizing free-viewpoint videos of human performances from sparse multi-view captures with generalization across subjects [7, 71, 10]. Neural Human Performer (NHP) [18] integrates pixel-aligned visual features with skeletal pose information extracted from multi-view video sequences through cross-attention. TransHuman [34] addresses performance degradation of NHP in occluded regions by learning to synthesize a NeRF-based representation in the canonical pose space with a transformer architecture that processes individual body parts as tokens. Neural Novel Actor (NNA) [57] disentangles appearance and geometry by processing spatial point features and SMPL surface features through a graph CNN, enabling separate prediction of person-specific appearance and pose-dependent deformation, as proposed by Liu et al. [25].
More recently, GPS-Gaussian [69] achieved real-time novel view synthesis of human performances by predicting per-pixel 3D Gaussians from 2 source views, which are unprojected to 3D space and aggregated for final rendering. Notably, GPS-Gaussian allows for view interpolation when source views have at most a 60-degree angular difference, thus requiring at least 6 input views in total. Generalizable Human Gaussians (GHG) [19] introduces an alternative approach for reconstructing static human models from sparse input views. By leveraging a large-scale dataset of textured human 3D scans [66], GHG demonstrates generalization to unseen subjects by representing virtual humans as multiple scaffolds of 3D Gaussians in the UV space. The most recent Vid2Avatar-pro [5] achieved animatable avatar generation from a monocular video by training on a large corpus of multi-view data. However, Vid2Avatar-pro requires a high-quality personalized body template at test time, which can only be reconstructed in a few hours for each novel identity.
Generative Methods. Another line of work focuses on single-image to animatable avatar reconstruction through generative modeling. The most common technique here is to rely on SMPL-X and its UV mapping to anchor 3D Gaussians. IDOL [72] employs a transformer-based encoder to combine image features with a learnable UV token to later decode 3D Gaussians with a convolutional network. Following IDOL, SIGMAN [64] learns a variational autoencoder [16] to map sparse views to 3D Gaussians, which is followed up by a diffusion transformer [37] to generate latents for novel avatars, conditioned on an image or a text prompt. Both IDOL and SIGMAN combine real-world datasets with synthetic data, generated with custom pipelines. MoGa [74] jointly optimizes a diffusion model to learn per-avatar latents and an autodecoder to convert latents to 3D avatars. At test time, MoGa optimizes a latent code for the input image and uses a multi-view diffusion model to generate additional pseudo ground truth views for this latent inversion process.
@@ 3 Method
In this section, we first formalize the problem setting and relevant background knowledge (Sec. 3.1). Then we introduce our generalizable human representation (Sec. 3.2), our training strategy (Sec. 3.3), and finally the implementation details (Sec. 3.4).
@@ 3.1 Problem Setting and Background
Problem Setting. GIGA aims at mapping single/sparse image observations (1-4 views) of an unseen human to a 3D Gaussian representation that can be rendered from a free viewpoint, faithfully preserving the appearance and clothing dynamics from these driving videos, even for unseen identities. Fig. 2 presents an overview of our method. To render dynamic sequences, GIGA processes each frame independently. Following the common settings [41, 48, 19], we assume that a motion tracking and a parametric body template are given to roughly describe the body shape and the pose of the human. Importantly, our method requires no personalized training.
For training, GIGA assumes a collection of multi-view videos of several hundred subjects with per-frame subject segmentation masks. Each subject is captured by at least $\hat{K}=16$ calibrated cameras, with $\pi_{\hat{k}}$ denoting the projection matrix for camera $\hat{k}$. Each video frame is annotated with SMPL-X [36] parameters released as additional labels in most multi-view human performance capture datasets  [61, 3, 66]. During training, we randomly sample 3 or 4 input views and use estimated SMPL-X parameters to construct the input for the encoder network. Images from the remaining views serve as ground truth to supervise the outputs decoded by the network.
@@ 3.2 Generalizable Human Representation
Inspired by previous personalized, animatable Gaussian avatar methods [35, 14, 56], we represent virtual humans as texel-aligned 3D Gaussian maps within the texture space $\mathcal{M}_{\mathrm{uv}}$ of the SMPL-X template mesh, displacing each 3D Gaussian in the canonical pose with a motion-dependent offset for better geometry modeling. Texel-aligned 3D Gaussians are semantically similar across various characters and 2D textures can be efficiently processed by convolutional neural networks. Thus, the predicted Gaussians from the same region of the texture space will be entangled thanks to the locality bias of convolutions. As such, GIGA can learn cross-subject information and predict the final texel-aligned 3D Gaussians with a single pass of a 2D convolutional neural network.
Input Appearance Encoding. We use information from the input views to capture identity-specific appearance with pose-dependent variations. To transform this information from image to texture space, we first compute partial textures $\mathbf{T}_{\mathrm{uv},k}$ for each view $\pi_{k}$ by projecting each pixel onto the 3D surface of SMPL-X. This 3D point can be mapped into texture space using the UV mapping of SMPL-X such that the final image pixel color is projected onto the respective 2D location in the texture map. We then fuse those textures according to visibility, resulting in the final texture $\mathbf{T}_{\mathrm{uv}}\in\mathbb{R}^{T\times T\times 3},T=512$ in the UV space $\mathcal{M}_{\mathrm{uv}}$.
@@
which extracts appearance features $\mathbf{F}_{\mathrm{uv}}^{a}\in\mathbb{R}^{T_{f}\times T_{f}\times d}$, that encode character-specific appearance and identity information. The encoder $\mathcal{E}_{\mathrm{a}}$ consists of 2D convolutional downsampling residual blocks. Feature maps $\mathbf{H}^{\mathrm{a}}$ from also taken from the downsampling levels for later usage in decoding. $\mathcal{E}_{\mathrm{a}}$ is conditioned on the motion embedding $\mathbf{y}_{\mathrm{m}}$. Following Rombach et al. [42], this conditioning is implemented using cross-attention blocks as final layers of the encoder.
Motion Embeddings. Since the input texture $\mathbf{T}_{\mathrm{uv}}$ contains only a localized pose-dependent learning signal from the template surface. We additionally use an MLP-based motion encoder $\mathcal{E}_{\mathrm{m}}$ to construct motion embeddings for SMPL-X poses $\boldsymbol{\theta}$, which adds global body motion awareness at the encoding stage:
@@
Gaussian Primitives Regression. For the decoding stage, we design three separate decoders (Fig. 2): $\mathcal{D}_{\mathrm{a}}$ for the appearance, $\mathcal{D}_{\mathrm{p}}$ for scales, quaternions and opacities of 3D Gaussians, and $\mathcal{D}_{\mathrm{g}}$ for per-Gaussian offsets. All three share the same convolutional architecture and receive appearance and geometry features $\mathbf{F}_{\mathrm{uv}}^{\mathrm{a}},\mathbf{F}_{\mathrm{uv}}^{\mathrm{g}}$ as inputs:
@@
The output map $\mathcal{G}^{\prime}_{\mathrm{uv}}$ contains RGB color channels $\mathbf{c}$, quaternions $\mathbf{q}_{0}$, normalized scales $\mathbf{s^{\prime}}$, opacities $\alpha$, and offsets $\delta\mathbf{x}_{0}$, which are defined in the canonical T-pose. To utilize the representational power of the shared texel space and to propagate semantic information from encoders to decoders at different spatial scales, we use UNet-like [44] skip-connections, building a feature propagation bridge from each encoder to each decoder. We stack intermediate feature maps $\mathbf{H}^{\mathrm{a}}$ and $\mathbf{H}^{\mathrm{g}}$ along the feature dimension and feed them to every decoder at each corresponding upsampling layer. To convert a texel-aligned 3D Gaussian map to the observed pose space $\boldsymbol{\theta}$, we use linear blend skinning applied to both offsets $\delta\mathbf{x}_{0}$ and quaternions $\mathbf{q}_{0}$: $\delta\mathbf{x}=\operatorname{lbs}\left(\delta\mathbf{x}_{0},\boldsymbol{\theta}\right);\mathbf{q}=\operatorname{lbs}\left(\mathbf{q}_{0},\boldsymbol{\theta}\right)$. We also multiply normalized Gaussian scales $\mathbf{s}^{\prime}$ by a hyperparameter $\rho$: $\mathbf{s}=\rho\mathbf{s}^{\prime}$, value $\rho=5e^{-3}$ was empirically found adequate to provide sufficient coverage of the template with Gaussians.